OpenStudy (anonymous):

inverse of the equation f(x)=1/3x+2 i think its g^-1(x)=3x-6? please help

OpenStudy (bibby):

looks good

OpenStudy (anonymous):

g^-1(x)=3x-6 that is correct?

OpenStudy (anonymous):

I just wanted to a make sure because when I graphed it, it didnt look like it would be opposite

OpenStudy (anonymous):

???

OpenStudy (daniel.ohearn1):

What makes you think it's that?

OpenStudy (daniel.ohearn1):

kathline?

OpenStudy (sloppycanada):

Yes, you're correct. That's what I got.

OpenStudy (bibby):

notice how any 2 points on the line are "reversed" between the line and its inverse. the easiest 2 to look at are the x and y intercepts

OpenStudy (anonymous):

because you first replace fx with y so y=1/3x-2, then solve for x interms of y so 1/3x=y-2 and get x=3y-6 then swap postions of variable y=3x-6 and then replace y with g(x) so therfore g^-1(x)=3x-6?

OpenStudy (anonymous):

awesome thank you so much!!